# ChemBOMAS: Accelerated BO in Chemistry with LLM-Enhanced Multi-Agent System

### v1

@@ Abstract
The efficiency of Bayesian optimization (BO) in chemistry is often hindered by sparse experimental data and complex reaction mechanisms. To overcome these limitations, we introduce ChemBOMAS, a new framework named LLM-Enhanced Multi-Agent System for accelerating BO in chemistry. ChemBOMAS’s optimization process is enhanced by LLMs and synergistically employs two strategies: knowledge-driven coarse-grained optimization and data-driven fine-grained optimization. First, in the knowledge-driven coarse-grained optimization stage, LLMs intelligently decompose the vast search space by reasoning over existing chemical knowledge to identify promising candidate regions. Subsequently, in the data-driven fine-grained optimization stage, LLMs enhance the BO process within these candidate regions by generating pseudo-data points, thereby improving data utilization efficiency and accelerating convergence. Benchmark evaluations further confirm that ChemBOMAS significantly enhances optimization effectiveness and efficiency compared to various BO algorithms. Importantly, the practical utility of ChemBOMAS was validated through wet-lab experiments conducted under pharmaceutical industry protocols, targeting conditional optimization for a previously unreported and challenging chemical reaction. In the wet experiment, ChemBOMAS achieved an optimal objective value of 96%. This was substantially higher than the 15% achieved by domain experts. This real-world success, together with strong performance on benchmark evaluations, highlights ChemBOMAS as a powerful tool to accelerate chemical discovery.
@@ 1 Introduction
Manual experimentation and traditional control variable methods have long underpinned chemical discovery, yet they remain labor-intensive and time-consuming, slowing the generation of new scientific insights Xie et al. [2023], Tom et al. [2024]. To address these constraints, automated or self-driving laboratories integrate robotic execution with AI algorithms, delivering high throughput, precision, and efficiency Seifrid et al. [2022], Chen et al. [2024], Ai et al. [2024a]. Within these experimental platforms, Bayesian Optimization (BO) algorithms are widely recognized as a crucial decision-making tool for experiment design Guo et al. [2023], Abolhasani and Kumacheva [2023], Chen et al. [2023], Ai et al. [2024b]. BO enables efficient navigation of complex experimental variable spaces and converges toward optimal reaction conditions or material compositions by integrating prior data, constructing probabilistic surrogate models, quantifying uncertainty, and iteratively selecting the most informative subsequent experimentsShields et al. [2021].
@@
Despite BO achieving remarkable success in complex scientific domains, particularly chemistry, it still contends with two major obstacles: (I) the scarcity and high cost of experimental observations during the early optimization stages, and (II) the multitude of reaction parameters that inflate the search into high-dimensional design spaces Shahriari et al. [2015], Wang et al. [2023]. The two obstacles exacerbate the limitations of vanilla BO, such as the "cold start" problem and the "curse of dimensionality", frequently leading to slow convergence Guo et al. [2023]. Without effective acceleration strategies, the protracted optimization process may yield only marginal improvements, which could cause researchers to abandon the search before discovering the optimal conditions.
@@
Several strategies have been proposed to accelerate BO, including search space partitioning Wang et al. [2020], specialized encoding embeddings Tripp et al. [2020], Moriconi et al. [2020], Nayebi et al. [2019], pseudo-data generation Yin et al. [2023], and transfer across similar tasks Swersky et al. [2013]. However, when these acceleration strategies are applied to the intricate chemical reactions, two critical shortcomings emerge. First, most approaches employ a single acceleration technique, which may prove insufficient for the chemical optimization problems with multiple demands, such as exploration of diverse reaction parameters combinations while enriching sparse early-stage data. Second, current acceleration methods are predominantly data-driven. Because reaction pathways differ widely in their underlying kinetics and thermodynamics, a purely statistical BO framework frequently expends resources in chemically implausible regions of the search space, missing opportunities to leverage mechanistic insight that could guide the search more efficiently.
To overcome these limitations, we propose ChemBOMAS, an LLM-Enhanced Multi-Agent System specifically designed for accelerated Bayesian Optimization in chemistry. ChemBOMAS introduces a novel approach where Large Language Models (LLMs) enhance a synergistic optimization process that combines knowledge-driven coarse-grained optimization with data-driven fine-grained optimization. The system operates using a multi-agent framework where agents collaborate dynamically. In the coarse-grained optimization stage, an LLM-powered agent reasons over existing chemical knowledge (e.g., literature, databases) to intelligently decompose the vast search space and identify promising candidate regions, effectively pruning chemically irrelevant areas. Subsequently, in the fine-grained optimization stage within these candidate regions, another LLM-enhanced workflow generates informative pseudo-data points. These pseudo-points augment the data available to the BO agent, significantly improving data utilization efficiency and accelerating convergence towards the optimum.
The effectiveness of ChemBOMAS was rigorously evaluated. We conducted extensive experiments on four benchmark chemical reaction datasets (Suzuki coupling, Buchwald-Hartwig coupling, Tandem arylation, and Enantio selectivity), demonstrating consistent improvements in optimization efficiency and results compared to various baseline methods. Furthermore, ablation studies confirmed the contribution of each component within our multi-agent framework. Crucially, the practical utility and real-world applicability of ChemBOMAS were validated through wet-lab experiments conducted under pharmaceutical industry protocols. These experiments targeted conditional optimization for a previously unreported and challenging chemical reaction, where ChemBOMAS achieved outstanding results, significantly surpassing the performance achieved by domain experts. Our main contributions are summarized as follows:
1. We propose ChemBOMAS, a novel LLM-enhanced multi-agent framework that synergistically integrates knowledge-driven and data-driven strategies for efficient chemical optimization, enabling dynamic collaboration between specialized agents.
2. We introduce an LLM-powered knowledge integration approach for coarse-grained optimization, where LLMs analyze domain knowledge to guide search space decomposition and identify high-potential regions.
3. We propose a fine-grained optimization stage in which an LLM-enhanced workflow synthesizes chemically plausible pseudo-experiments within the high-potential regions, augmenting the sparse experimental observations and accelerating convergence toward the global optimum.
4. We demonstrate the effectiveness and efficiency of ChemBOMAS through comprehensive experiments on diverse benchmark datasets and, critically, through successful wet-lab validation on a novel and challenging chemical reaction, outperforming domain experts.
@@ 2 Related Work
Large Language Models (LLMs) offer synergistic potential with Bayesian Optimization (BO) to address traditional BO limitations (e.g., sample inefficiency, cold starts) by providing prior knowledge Souza et al. [2021], enhancing surrogate models Liu et al. [2024], Nguyen et al. [2024], Ramos et al. [2023], automating acquisition function design Austin et al. [2024], and enabling optimization in novel problem representations. Prior work has explored LLM-driven BO improvements in warm-starting, surrogate modeling, candidate generation, acquisition function design, and search space understanding.
@@
However, directly replacing core BO modules with LLMs introduces significant challenges. LLM "hallucinations" can mislead optimization, compromising reliability. Furthermore, the direct suitability of LLMs as surrogates or acquisition functions is limited by concerns regarding uncertainty quantification, theoretical guarantees, computational cost, efficiency in low-data regimes, adaptability to specific numerical tasks, and interpretability. On another front, some techniques such as LA-MCTS Wang et al. [2020] was proposed, which employ tree structures to decompose the search space Wang et al. [2023, 2024a, 2019]. Some works propose hierarchical Bayesian optimization Moriconi et al. [2020], Reker et al. [2020]. These approaches offer valuable strategies for managing and navigating complex optimization landscapes. Unlike previous works, we focus on robustly integrating LLM knowledge to enhance BO, leveraging their strengths as auxiliary tools while mitigating weaknesses such as hallucinations, to achieve this synergy over substitution.
3 Background
The discovery and optimization of chemical reactions are pivotal for progress in diverse fields, including pharmaceuticals, materials science, and sustainable energy solutions. Fundamentally, this endeavor can be framed as a classic black-box optimization problem. The objective is to identify an optimal set of reaction parameters from a multi-dimensional search space that maximizes a desired reaction performance, which is quantified by an objective function . defines the multi-dimensional search space and represent the spaces of experimental conditions such as temperature, pressure, catalyst type, solvent, reactant concentrations, and reaction time. Sometimes it also involves the selection or modification of reactants (e.g., substituting a functional group). Desired reaction performance is typically quantified by metrics such as yield (the amount of desired product obtained from a chemical reaction), selectivity (the ratio of desired product to undesired byproducts), or a composite objective function combining multiple criteria. Formally, this optimization problem is expressed as finding , where represents the reaction parameters and the objective function maps these parameters to the experimental performance.
4 ChemoBOMAS
To address the identified challenges, we present ChemBOMAS, a pioneering two-stage multi-agent optimization framework designed to enhance the efficiency of reaction condition optimization. As depicted in Figure 1, the framework is systematically structured into two distinct stages: coarse optimization through knowledge-driven search space decomposition and data-driven fine-grained optimization. By seamlessly integrating these phases, ChemBOMAS offers a comprehensive and effective approach to navigating the complexities inherent in chemical reaction optimization. The following sections provide a comprehensive exposition of the two stages, detailing their methodologies and underlying principles.
4.1 Knowledge-driven search space decomposition
This stage leverages the literature comprehension and reasoning capabilities of large language models to systematically analyze relevant literature and physicochemical data, assessing the significance of various reaction parameters on the outcome and identifying key physicochemical properties that influence reaction performance. Based on the importance of reaction parameters and the corresponding physicochemical characteristics, a hierarchical tree structure is employed to partition the condition space, facilitating a coarse-grained search for optimal reaction conditions and enhancing the efficiency of experimental exploration.
Specifically, the large language model (LLM) first conducts a systematic search of chemical literature relevant to the target reaction, extracting and synthesizing key information to rank the potential significance of various reaction parameters, including reactants, catalysts, ligands, solvents, and bases. Furthermore, the LLM identifies critical physicochemical properties, such as electronic and steric effects, that influence the experimental outcomes. Subsequently, it retrieves specific physicochemical properties for all conditions within the reaction space from a physicochemical property database and performs clustering based on these properties.
A hierarchical optimization tree is then constructed based on the importance ranking of reaction parameters and the clustering results based on their physicochemical property. The root node of this tree represents the complete parameter space (All possible combinations of reaction parameters) for the reaction to be optimized and the other nodes of the tree are the subsets of the whole parameter space. The hierarchical structure of the tree is built layer by layer according to the importance ranking of the parameters (the more important condition variable corresponds to the layer with a higher level). At each layer, the parent node from the previous layer is split into several child nodes based on the physicochemical property of the corresponding condition variable of this layer. This splitting process is performed recursively until all considered parameters are incorporated into the tree structure, ultimately forming a decision tree that guides the optimization direction. To facilitate a deeper understanding of the entire process, a detailed illustrative example is provided in the supplementary materials for further clarification.
Once the optimization tree is constructed, it can be systematically traversed layer by layer to identify and sample high-value subspaces of reaction conditions. It facilitates the reduction of the optimization search space, thereby enhancing efficiency in the refinement process. Specifically, we employ the Upper Confidence Bound (UCB) algorithm to search this tree. Starting from the root node, we iteratively select the subsequent child nodes at each hierarchical level based on the UCB strategy, conducting in-depth exploration until reaching the final leaf node. For each child node i of the current parent node, we calculate its UCB value: , where is the average experimental return (e.g., average yield or selectivity) of child node i, representing the ‘exploitation’ term. It is calculated by dividing the total return obtained from this node by its number of visits . is the total number of times the current parent node has been visited. is the number of times child node i has been visited. is an exploration constant used to balance the relative importance of the exploitation and exploration terms.
At each hierarchical level, the child node exhibiting the highest UCB value is selected for the subsequent exploration step. This formula ensures that while the algorithm tends to select child nodes that have performed well historically, it also gives an opportunity for exploration to less-visited child nodes, thus avoiding premature convergence to a suboptimal solution.
As the selection process recursively traverses a path within the tree until a leaf node is reached, the final subset of chemical reaction conditions is established. Subsequently, Bayesian optimization is adopted to sample reaction conditions and corresponding experiments (or, in simulations, query results in datasets) are conducted to evaluate the actual impact of these reaction conditions, obtaining key performance indicators such as reaction yield and selectivity. This evaluation value, denoted as , serves as a reward signal. This signal is then backpropagated to update the statistics of all nodes along the path from the currently evaluated leaf node to the root node. The specific update rules are as follows:
1. The visit count for each node on the path is updated by .
2. The cumulative return for each node on the path is increased by : .
3. Through these updates, the average return of the node is also adjusted accordingly, thereby influencing the UCB value calculation and node selection in subsequent iterations.
The entire process consists of three key steps: (1) trajectory sampling on the tree to obtain the final subset of reaction conditions, (2) Bayesian optimization within this subset to sample reaction conditions and conduct experiments to acquire the reward signal, and (3) backpropagation to update the Upper Confidence Bound (UCB) values of the nodes. Through continuous iterations of these steps, the value estimation of nodes or chemical condition subsets becomes increasingly precise, thereby enhancing the guidance for optimal condition selection.
4.2 Data-driven Bayesian Optimization
The knowledge-guided optimization tree constructed can effectively decompose the search space and capture potential correlations under specific conditions (e.g., certain ligands are not used with strong bases). However, relying solely on knowledge makes it difficult to accurately predict reaction performance, and obtaining specific reaction variable recommendations still requires intrinsic information of experimental data. Given the high cost and lengthy cycles of real experiments, as well as the cold-start problem of traditional Bayesian Optimization (BO), the Large Language Model (LLM) is used to learn from historical experimental data and mine underlying complex patterns to provide a higher quality prior for BO’s surrogate model, thereby accelerating the optimization process. The LLM exhibit good compatibility in handling data with different representations, can be pre-trained using massive amounts of unlabeled data and non-domain-specific data, and support diverse task types. Particularly noteworthy is their potential for extrapolation, which offers possibilities for exploring unknown reaction spaces. Therefore, we apply LLMs to the task of predicting the reaction performance.
To achieve reaction performance prediction, we first construct an LLM-based regression model. The model directly takes SMILES (Simplified Molecular Input Line Entry Specification) strings as input. SMILES is a linear notation that encodes the topological structures of molecules, including atoms, chemical bonds, rings, and branches, into one-dimensional ASCII text, facilitating efficient computer processing, storage, and widespread application in structure retrieval and analysis within chemoinformatics and molecular modeling research. We selected LLaMa 3.1 Grattafiori et al. [2024] as the base architecture.
We employ the Pistachio dataset for pre-training. The Pistachio dataset (a proprietary dataset acquired from a commercial vendor specializing in chemical data) primarily integrates various patent data (e.g., from USPTO). Considering that the Pistachio dataset may suffer from issues such as missing entries, difficulty in tracing sources, and significant noise in key performance data like reaction yields, we do not directly incorporate reaction performance information during the pre-training phase. Instead, we design a conditional prediction task: given the reactants and products (i.e., the reaction equation), the model’s objective is to predict the corresponding reaction conditions. This process utilizes a Causal Language Modeling (CLM) loss, which involves predicting the next token in the sequence of reaction conditions: , where denotes the token index, represents the -th token in the condition sequence, and is the sequence length. To enhance chemical awareness, we augment input sequences with functional group annotations using the RDKit toolkit Landrum et al. [2019], enabling explicit modeling of reaction-critical substructures.
Given that our research primarily focuses on predicting continuous variables like yield and selectivity, we enhance a pre-trained LLM by integrating a regression head. Through prompt engineering, we guide the LLM, which takes an input (e.g., prompted reaction conditions), to generate token sequences pertinent to reaction performance. The hidden state of the last token from this LLM output, which we can denote as after processing the input, is then fed into an additionally incorporated projection layer. This layer, for which we adopted a Multilayer Perceptron (MLP), , outputs the specific reaction performance prediction, aiming for the ground truth value .
During the subsequent fine-tuning stage, we optimize the entire architecture using a small amount of labeled data, , consisting of pairs. This process employs two distinct parameter update strategies:
The pre-trained LLM component is fine-tuned using Low-Rank Adaptation (LoRA) technology Hu et al. [2022]. This introduces a set of low-rank adaptable parameters, , to the LLM’s original weights. This approach effectively alleviates catastrophic forgetting Li et al. [2025], ensuring the LLM retains pre-trained knowledge while adapting its representations, now denoted , to the new task.
The MLP, now more precisely , is trained with its full set of parameters to accurately map the final token representation from the adapted LLM to the target output .
The overall fine-tuning objective is to minimize an L2 loss function with an L2 regularization term on the MLP’s parameters controlled by a hyperparameter . The complete fine-tuning loss is:
This formulation jointly optimizes the LoRA parameters of the LLM and the full parameters of the MLP to enhance predictive accuracy on target variables like yield and selectivity.
The trained LLM prediction model can be used to generate pseudo-points, providing data prior for BO’s surrogate model. By performing inference over the entire search space (or the decomposed subspaces), we can obtain pseudo-labels (i.e., predicted reaction performance) for all combinations of reaction parameters. These pseudo-points are then used to initialize the BO surrogate model. Although this step involves a considerable amount of inference computation, it is still negligible compared to the high costs and time investment of chemical experiments. Notably, performing BO in decomposed subspaces also helps reduce overall computational complexity. As BO iterates and real experimental data accumulates, we will gradually remove pseudo-points to increase the weight of real data in model updates, thereby mitigating the potential impact of pseudo-data noise. The removal of pseudo-points follows two core principles:
Data Similarity (Local Removal): We utilize the embedding of the LLM’s last token to calculate the cosine similarity between data points. If the similarity between a pseudo-point and a newly acquired real feedback point (i.e., actual experimental result) exceeds a predefined threshold, these pseudo-points are removed from the pseudo-dataset.
Observed Performance (Global Removal): Considering that as iteration progresses, the observed optimal performance tends to improve, the model should be encouraged to explore more broadly rather than over-relying on existing predictions. Therefore, based on the predicted performance values of the pseudo-points, we randomly discard a certain proportion of pseudo-data, starting from those with high predicted performance downwards.
These generated pseudo-points can also provide further support for the construction of the knowledge-guided optimization tree. By adjusting the temperature parameter of the LLM during prediction generation, we can guide the model to produce different rankings and classifications of reaction condition combinations’ performance, thereby generating multiple candidate optimization tree structures. Using a large number of pseudo-points, we can quantitatively evaluate the rationality of these candidate trees. Ideally, reaction varibales combinations within the same tree node should exhibit similar performance characteristics. Therefore, we calculate the variance of the predicted performance of all pseudo-points within each node for every candidate tree. The tree structure that minimizes the sum (or average) of intra-node variances is selected as the optimal knowledge-guided optimization tree.

### v2

# ChemBOMAS: Accelerated BO for Scientific Discovery in Chemistry with LLM-Enhanced Multi-Agent System

@@ Abstract
Bayesian optimization (BO) is a powerful tool for scientific discovery in chemistry, yet its efficiency is often hampered by the sparse experimental data and vast search space. Here, we introduce ChemBOMAS: a large language model (LLM)-enhanced multi-agent system that accelerates BO through synergistic data- and knowledge-driven strategies. Firstly, the data-driven strategy involves an 8B-scale LLM regressor fine-tuned on a mere 1% labeled samples for pseudo-data generation, robustly initializing the optimization process. Secondly, the knowledge-driven strategy employs a hybrid Retrieval-Augmented Generation approach to guide LLM in dividing the search space while mitigating LLM hallucinations. An Upper Confidence Bound algorithm then identifies high-potential subspaces within this established partition. Across the LLM-refined subspaces and supported by LLM-generated data, BO achieves the improvement of effectiveness and efficiency. Comprehensive evaluations across multiple chemical benchmarks demonstrate that ChemBOMAS set a new state-of-the-art, accelerating optimization efficiency by up to 5-fold compared to baseline methods.
@@ 1 Introduction
Manual experimentation and traditional control variable methods have long underpinned chemical discovery, yet they remain labor-intensive and time-consuming, slowing the generation of new scientific insights xie2023toward, tom2024self. To address these constraints, automated or self-driving laboratories integrate robotic execution with AI algorithms, delivering high throughput, precision, and efficiency seifrid2022autonomous, chen2024robot, ai2024customizable. Within these experimental platforms, Bayesian Optimization (BO) algorithms are widely recognized as a crucial decision-making tool for experiment design guo2023bayesian, abolhasani2023rise, chen2023robot, ai2024demand. BO enables efficient navigation of complex experimental variable spaces and converges toward optimal reaction conditions or material compositions by integrating prior data, constructing probabilistic surrogate models, quantifying uncertainty, and iteratively selecting the most informative subsequent experiments Shields2021.
@@
Despite BO achieving remarkable success in complex scientific domains, particularly chemistry, it still contends with two major obstacles: (I) the scarcity and high cost of experimental observations during the early optimization stages, and (II) the multitude of reaction parameters that inflate the search into high-dimensional design spaces shahriari2015taking, wang2023recent. The two obstacles exacerbate the limitations of vanilla BO, also known as the "cold start" problem and the "curse of dimensionality", frequently leading to slow convergence guo2023bayesian. Without effective acceleration strategies, the protracted optimization process may yield only marginal improvements, which could cause researchers to abandon the search before discovering the optimal conditions.
@@
Several strategies have been proposed to accelerate BO, including search space partitioning wang2020, specialized encoding embeddings tripp2020sample, moriconi2020high, nayebi2019framework, pseudo-data generation yin2023high, and transfer across similar tasks swersky2013multi. However, when these acceleration strategies are applied to the intricate chemical reactions, two critical shortcomings emerge. First, most approaches employ a single acceleration technique, which might be insufficient for the chemical optimization problems with multiple demands, such as exploration of diverse reaction parameter combinations while overcoming data scarcity in the early-stage. Second, current acceleration methods are predominantly data-driven. Because chemical reaction pathways differ widely in their underlying kinetics and thermodynamics, a purely statistical BO framework frequently expends resources in chemically implausible regions of the search space, missing opportunities to leverage mechanistic insight that could guide the search more efficiently.
To overcome these limitations, we propose ChemBOMAS, an LLM-Enhanced Multi-Agent System specifically designed for accelerated Bayesian Optimization in chemistry. ChemBOMAS synergistically integrates two LLM-powered modules: a knowledge-driven search space decomposition module and a data-driven pseudo-data generation module. The knowledge-driven module employs an LLM-powered agent to reason over existing chemical knowledge (e.g., literature, databases), intelligently decompose the vast search space and identify promising candidate regions, dynamically pruning the search space for better BO efficiency. Simultaneously, the data-driven module utilizes a fine-tuned LLM to generate informative pseudo-data points across the entire search space. These pseudo-data not only warm-start the BO process but also inform the knowledge-driven module’s subspace selection. This closed-loop interaction enables ChemBOMAS to achieve superior optimization efficiency and convergence speed even under extreme data scarcity.
The effectiveness of ChemBOMAS was rigorously evaluated. We conducted extensive experiments on four chemical performance optimization benchmarks, demonstrating consistent improvements in optimal results, convergence speed, initialization performance, and robustness compared to various baseline methods. Crucially, ablation studies confirmed that the synergy between the knowledge-driven and data-driven strategies is essential for creating a highly efficient and robust optimization framework. Additionally, the practical utility and real-world applicability of ChemBOMAS were validated through a previously unreported wet-lab experiment.
Our main contributions are summarized as follows:
1. We systematically investigated how LLM-based approaches could address two key limitations in BO for scientific discovery in chemistry: data scarcity and inefficiency in the vast search spaces.
2. We propose ChemBOMAS, a novel framework that synergistically combines LLM-powered knowledge and data strategies, which enables efficient BO in chemical performance optimization tasks. It consists of a knowledge-driven module to decompose the search space and a data-driven module to generate pseudo-data.
3. ChemBOMAS significantly accelerates BO under limited data (1% of all search space) on four chemical benchmarks. It achieves accelerated convergence and superior final performance compared to four relevant baseline methods, improving optimal results by approximately 3–10% and converging about 2–5 faster.
@@ 2 Related Work
Large Language Models (LLMs) offer synergistic potential with Bayesian Optimization (BO) to address traditional BO limitations (e.g., sample inefficiency, cold starts) by providing prior knowledge DBLP:conf/pkdd/SouzaNOOLH21, enhancing surrogate models liu2024, DBLP:journals/corr/abs-2410-10190, ramos2023, automating acquisition function design DBLP:conf/recsys/AustinKTS24, and enabling optimization in novel problem representations. Prior work has explored LLM-driven BO improvements in warm-starting, surrogate modeling, candidate generation, acquisition function design, and search space understanding.
@@
However, directly replacing core BO modules with LLMs introduces significant challenges. LLM "hallucinations" can mislead optimization, compromising reliability. Furthermore, the direct suitability of LLMs as surrogates or acquisition functions is limited by concerns regarding uncertainty quantification, theoretical guarantees, computational cost, efficiency in low-data regimes, adaptability to specific numerical tasks, and interpretability.
On another front, some techniques such as LA-MCTS wang2020 was proposed, which employ tree structures to decompose the search space wang2023recent, wang2024, wang2019. Some works propose hierarchical Bayesian optimization moriconi2020high, reker2020adaptive. These approaches offer valuable strategies for managing and navigating complex optimization landscapes. Unlike previous works, we focus on robustly integrating LLM knowledge to enhance BO, leveraging their strengths as auxiliary tools while mitigating weaknesses such as hallucinations, to achieve this synergy over substitution.
3 Methodology
3.1 Problem Setup
This work aims to significantly improve the efficiency of searching a task’s variable space for the optimal combination that maximizes the objective function.
3.2 The Framework of ChemBOMAS
As illustrated in Figure 1, we propose ChemBOMAS, an LLM-enhanced multi-agent optimization framework that systematically integrates data-driven and knowledge-driven strategies. First, the data-driven strategy utilizes a pre-trained and fine-tuned LLM regressor to generate pseudo-data, thereby robustly initializing the optimization process. Second, the knowledge-driven strategy employs a hybrid Retrieval-Augmented Generation (RAG) approach, which guides an LLM to partition the search space based on variable impact ranking and property similarity. Third, an Upper Confidence Bound (UCB) algorithm then identifies the most promising subspaces from this partition. Finally, BO is performed within the selected subspaces, supported by the LLM-generated pseudo-data, leading to enhanced effectiveness and efficiency. The complete algorithm process can be seen in Appendix F.The two strategies are detailed below.
3.3 Data-driven Strategy: LLM-generated Pseudo Data
An LLM-based regression model was constructed and utilized in three sequential steps to generate pseudo-data for optimization initialization.
Step 1: Pre-training. The base LLaMA 3.1 model grattafiori2024llama3herdmodels was pre-trained on the Pistachio dataset to enhance its representational ability for chemical reactions. The dataset was formatted as Q&A pairs where, given reactants and products , the model learns to predict the corresponding reaction conditions , thereby avoiding direct exposure to objective performance labels. Pre-training employed a Causal Language Modeling loss: , where denotes the token index, represents the -th token, and is the sequence length.
Step 2: Fine-tuning. The pre-trained model was subsequently fine-tuned on a small labeled dataset , which only comprises 1% of the data, by integrating a regression head. A reaction configuration (including , , and ) is fed into the LLM via prompt engineering. The final hidden state is then projected to a reaction performance prediction via an MLP: . Fine-tuning used Low-Rank Adaptation (LoRA) DBLP:conf/iclr/HuSWALWWC22 with rank , introducing adaptable parameters alongside the frozen pre-trained weights . The MLP parameters were fully trained to minimize an L2-loss with regularization:
Step 3: Pseudo-data Generation and Utilization. The fine-tuned LLM regressor was used to generate pseudo-data for all unsampled data points, forming a pseudo-dataset , where was used to initialize a UCB algorithm. The UCB algorithm (Section 3.4) then identifies the high-potential subspaces. BO is then conducted within these subspaces (Section 3.5), leveraging both the selected pseudo-data and limited real data to accelerate surrogate model fitting. To mitigate the influence of the noise in the pseudo-data, a refinement strategy based on data similarity and reverse-order removal is applied (see Appendix D for details).
3.4 Knowledge-driven Strategy: LLM-Divided Search Space
To efficiently identify high-potential regions in the vast chemical reaction parameter space, we construct a Subspace Tree Search module using the GPT-4o API in three steps.
Step 1: LLM-Guided Space Partitioning. The -dimensional variable space is defined as , containing categories of chemicals involved in the reaction, where each represented a category including candidate substances. A hybrid Retrieval-Augmented Generation (RAG) approach integrates multi-source information (literature, professional databases, web search) to facilitate the LLM’s decisions and minimize hallucination. The LLM first ranks the chemical categories by their importance to the chemical reaction, generating an ordered sequence . Subsequently, for each chemical category , the LLM identifies key influencing physicochemical properties and clusters the candidates based on these property values. This partitions each into a collection of seperate subspaces , where candidate substances within each subspace share similar properties.
Step 2: Hierarchical Search Tree Construction. A hierarchical tree is built based on the category importance order and the clustering results . The -th layer of the tree corresponds to the -th most important category and contains nodes representing its clusters. Each path from the root to a leaf node defines a unique search subspace as the Cartesian product of clusters, resulting in a total of disjoint subspaces that comprehensively partition the original space.
Step 3: UCB-based Subspace Selection. A UCB algorithm is employed to explore the tree and identify promising subspaces. Starting from the root, UCB selects child nodes layer-by-layer until reaching a leaf node. The UCB value for a child node is computed as: , where is the average performance value (exploitation), is the parent’s visit count, is the child’s visit count, and is an exploration constant. At each layer, the top-5 nodes by UCB value are selected for further exploration. This path traversal pinpoints high-value subspaces for subsequent BO. The UCB values are updated dynamically as BO progresses and new samples are acquired.
3.5 Bayesian Optimization in ChemBOMAS
BO is performed within the promising subspaces identified by the preceding modules, leveraging the LLM-generated pseudo-data for initialization. The procedure consists of two main steps.
Step 1: Surrogate Modeling. A Gaussian Process (GP) surrogate model, using a Matérn kernel with constant scaling and a white noise kernel, is fitted to the combined set of actual observations and pseudo-data points, which serve as an informative prior. This model provides, for any unsampled point in the target subspaces, a posterior distribution over the performance value, characterized by a mean function and a variance function .
Step 2: Acquisition Function Optimization. An acquisition function , such as UCB or Expected Improvement (EI), is used to recommend the next sample by balancing exploration (high uncertainty) and exploitation (high predicted mean). The next query point is selected by maximizing over the unsampled points within the high-potential subspaces: . This point is then evaluated to obtain a new real observation, which updates the GP surrogate model for the next iteration.